OpenStudy (anonymous):

25.06/25.02=a^(.03) How do I convert the exponent. If the exponent was a 3 I know I could multiply both sides by (1/3). But this decimal is messing me up. haha Thanks!

OpenStudy (zzr0ck3r):

if the decemel was 3 you could multiply both sides by 1/3 ?

OpenStudy (zzr0ck3r):

you can solve this with logs

OpenStudy (anonymous):

Can also convert 0.03 to the fraction 3/100 and use exponent rules.

OpenStudy (anonymous):

zzRock If the Exponent was 3 We aren't up to the logs section yet thats why I don't want to use logs.

OpenStudy (anonymous):

I'd be curious to see from zzrocker how to do this with logs since we're solving for the base and not the exponent.

OpenStudy (anonymous):

My reasoning is that if you wanted to solve \[A=x^2 \rightarrow x=A^{1/2}\] Likewise, if \[A=x^{1/2} \rightarrow x=A^2\] So, if \[K=a^{3/100} \rightarrow a=K^{100/3}\]

OpenStudy (anonymous):

(This is all assuming no negative numbers and no complex numbers)

OpenStudy (anonymous):

So I'm sure when you said, " If the exponent was a 3 I know I could multiply both sides by (1/3)" you meant that you could raise both sides to the exponent 1/3. That is correct. Here, the exponent is 3/100, so exponentiate both sides to the 100/3.

OpenStudy (anonymous):

OK thanks Cliff

OpenStudy (anonymous):

And yes raise to the power of (1/3) in my example :)

OpenStudy (anonymous):

Let me know what you get as your answer. I got something close to 1.05

OpenStudy (anonymous):

Yes I got 1.055. Now I am supposed to use this to figure out the growth rate. So f(.05)=25.06 I will be using Q=Q0a^t 25.06=Q0(1.055)^.05 So when I do (1.055)^.05 I get 1.003 Then I 25.06/1.003 to get 24.993. But thats supposed to be percentage growth rate so 2499%?!

OpenStudy (anonymous):

Not sure if I know what you mean. A percent increase from 25.02 to 25.06, or do you want the growth factor? And what does that .05 mean?

OpenStudy (anonymous):

t would be x P is an exponential function of t with base a

OpenStudy (anonymous):

Ah, ok, so t=0.05 is like .05 seconds or .05 days or something like that?

OpenStudy (anonymous):

Ok, I think I followed your above steps. First of all, Is your function f(t)=Q0*a^t or Q0*a^(kt)? Secondly, if you divide 25.06 by (1.055)^0.05 you are solving for Q0, right?

OpenStudy (anonymous):

If you're looking for percentage increase then it depends on how you want to express that. You can say that 25.06 is 25.06/Q0 times as large as Q0, or you can say the amount of increase (25.06 - Q0) is [(25.06 - Q0)/Q0]*100% more than Q0..

OpenStudy (anonymous):

here is the original problem f(.02)=25.02 and f(.05)=25.06

OpenStudy (anonymous):

I just want to make sure I have any idea what is going on maybe I am screwing it up somewhere

OpenStudy (anonymous):

I see. f(.02)=25.02 and f(.05)=25.06 was the given information and the question is "what is the growth rate?"

OpenStudy (anonymous):

Then I think you should probably start over.

OpenStudy (anonymous):

You missed the growth rate variable from your original equation. I think that is what you are supposed to solve for. You may need logarithms afterall.

OpenStudy (anonymous):

does a^.03 change into 1/.03 to get rid of exponent? becuase a^3 would be 1/3 to get rid of the exponent

OpenStudy (anonymous):

Neither. The equation you need to start with is \[f(t)=P_0e^{rt}\] Your task is to solve for r.

OpenStudy (anonymous):

I hope I'm understanding your question right because this will require logarithms, and you said that you didn't want to use logs.

OpenStudy (anonymous):

They are using P=P0a^t in the book I can do the problem if it isn't a decimal for t! lol

OpenStudy (anonymous):

Ok, so it's replacing e^r with a, I see..

OpenStudy (anonymous):

Hmm, then what you did is fine, that gets you a, but then I'm not sure what it means by find growth rate if you're not finding r.

OpenStudy (anonymous):

ok thanks Ill get help on it tomorrow

OpenStudy (anonymous):

1.5 hours is enough on one problem when I can do the others so easily its really irritating

OpenStudy (anonymous):

Ok, you can still get the growth rate, r from here, but you still need to use logarithms. Understood. Leave it alone for now and reread it with a fresh mind.